# Security of AI Agents

### v2

@@ Abstract
The study and development of AI agents have been boosted by large language models. AI agents can function as intelligent assistants and complete tasks on behalf of their users with access to tools and the ability to execute commands in their environments, Through studying and experiencing the workflow of typical AI agents, we have raised several concerns regarding their security. These potential vulnerabilities are not addressed by the frameworks used to build the agents, nor by research aimed at improving the agents. In this paper, we identify and describe these vulnerabilities in detail from a system security perspective, emphasizing their causes and severe effects. Furthermore, we introduce defense mechanisms corresponding to each vulnerability with meticulous design and experiments to evaluate their viability. Altogether, this paper contextualizes the security issues in the current development of AI agents and delineates methods to make AI agents safer and more reliable.
1 Introduction
AI agents are robots in cyberspace, executing tasks on behalf of their users. To understand their user’s command, they send the input prompts as requests to foundation AI models, such as large language models (LLMs). The responses generated by the model may contain the final actions of the agent or further instructions. To execute the actions, the agent invokes tools, which may run local computations or send requests to remote hosts, such as querying search engines. The tools output results and feedback to the AI model for the next round of actions. By invoking tools, AI agents are granted the ability to interact with the real world. Since AI agents depend on their AI model to understand user input and environment feedback and generate actions to use tools, we say that the AI model is the backbone of the agent. We summarize the basic architecture of LLM-based AI agents in Figure 1. Traditional agents operate on pre-defined rules [castelfranchi1998modelling, wilkins2014practical] or reinforcement learning models [isbell2001social], making them hard to generalize to new tasks and different tools. LLM-based AI agents, on the contrary, can be practical in various tasks benefiting from enormous pre-training knowledge and the ability to read tool documentation as additional prompts. We use the term AI agent to denote all LLM-based agents in this paper.
Over the years, AI agents have showcased their outstanding performance on tasks including but not limited to writing shell scripts to interact with operating systems, querying databases, shopping and browsing on the web, and playing video games [yao2022webshop, park2023generative, zhou2023webarena, liu2023agentbench]. Agent-based personal assistants are also gaining popularity in society, for example, Microsoft [semantic_kernel, copilot] and HuggingFace [hfagents]. Overall, AI agents are a promising step forward toward the goal of robots with artificial general intelligence (AGI).
@@ 1 Introduction
Despite their popularity, existing research and development of AI agents failed to take into account their potential vulnerabilities. In traditional computing systems, security is guarded by three properties: confidentiality, integrity, and availability. Each of these properties faces unique challenges in AI agent usage.
Confidentiality is often managed by model-based access control policies, which abstract the system components and users into subjects, objects, and rights [bishop2004introduction]. However, these principles face significant challenges when applied to LLM-based systems due to the nature of LLMs to memorize [mireshghallah2022memorization, carlini2023quantifying, tirumala2022memorization] and compress [deletang2023language] training data. AI agents are granted the ability to interact with tool applications by reading their instructions and feedback, leaving more possibilities for privacy leaks. The ability to use tools introduces additional layers of complexity in maintaining confidentiality. As a result, we have to rethink information confidentiality in the context of AI agents. When assisting users with automatic tool usage, requests for sensitive information are unavoidable. This evaluation is essential to address the unique challenges posed by AI agents, especially when they are learning from user chat history and tool interaction logs, to ensure that data privacy protections evolve to effectively safeguard information in this new technological landscape.
@@
The threat of availability should be re-investigated for AI agents as well. Systems, data, and applications should always be available when the users need them. Unlike LLMs, which are stateless and can only output text tokens, AI agents execute actions that could affect the computing system itself. Therefore, each of the agent’s actions may have its own vulnerabilities to the agent’s host machine and tools. Current development on AI agents evaluates them in benchmark settings [liu2023agentbench, zhou2023webarena, park2023generative], failing to consider the difference between benchmark environments and real-world applications. AI agents without sanitization can harm the availability of both its host system and its tools by executing malicious commands generated by its LLM. To clarify between these vulnerabilities and the security of LLMs, malicious actions might be generated by hallucinations and prompts that do not break LLM’s alignment and prompt filter, requiring different defenses and safeguarding.
@@
In this paper, we explain and discuss the possible security issues of AI agents in detail. To facilitate future research, we propose several defense methodologies for the vulnerabilities we discovered on the component level in the AI agent architecture. To evaluate our defense proposals, we also set up preliminary experiments that our solutions depend on. Our contributions are as follows:
  • We formally introduce potential vulnerabilities of AI agents, and explain the causes and effects of these vulnerabilities in detail.
  • We propose multiple defenses to close the gap between AI research and AI agents in practice.
  • We verify the applicability of our proposed defenses and discuss their limitations and directions for future improvement.
2 Threat model
@@
We assume the AI agent is unimodal, where the input to the AI agent is text-only. We assume that the server that runs the AI agent is secure. Users can only access the server via the API provided by the AI agent. The programs that the AI agent runs have no undefined behavior, such as buffer overflow that allows remote code execution. We assume the AI agent has access to one or multiple tools, and will execute the tools solely based on the actions generated by the LLM.
Adversary’s capabilities
In our threat model, we assume the attackers are users of the agent and have the knowledge to attack a machine learning model. The attackers can perform adversarial examples/prompts, data poisoning, and network-based attacks on the AI agent using the provided interface. To be more precise, we assume the adversary will not directly attack the backbone tools and system of the AI agent. We assume the adversary can break the prompt guarding and alignment of the backbone LLM through interaction.
Difference between LLM and AI agent security
The threats for language models and AI agents are different. AI agents have different threats compared to language models. For generative models, one of the largest threats is jailbreaking, which aims to break the alignment of the model. After the pre-training stage, generative models usually experience alignment fine-tuning, including reinforcement learning from human feedback (RLHF) [askell2021general, bai2022training] and red-teaming [perez2022red], to improve the safety of the AI-generated content. However, rather than causing vulnerabilities in real-world computing systems, these issues are focused on the alignment between the model and human values, including ethics, offensive language, and politics. In contrast, since AI agents use LLMs to generate actions and have access to tools, they can be threats to real computing systems, applications, and resources, thereby introducing dangers to confidentiality, integrity, and availability in novel domains. In this paper, we focus on the potential threats caused by AI agents.
3 Potential vulnerabilities
In this section, we identify the important potential vulnerabilities that an AI agent application faces. Some of these vulnerabilities are akin to LLM vulnerabilities, but can happen through different means or require different defenses in the AI agent setting. The other vulnerabilities we find are specific to AI agents because of their ability to interact with the real world and execute actions generated by LLMs.
3.1 Sessions
@@
where $Q$ is the question queries and $A$ is the answers from the LLM. In simpler terms, we consider the language model to be “honest,” meaning it always generates the same response when given the same question. Therefore, the AI agent is responsible for managing the state of its LLM. If the AI agent has only one API account on the AI model, then instructing the AI model to separate the sessions of different users raises concerns on information leakage and action mis-assignment. On the other hand, even if the AI agent has multiple API accounts on the AI model, mapping user sessions to API accounts faces the same vulnerabilities when the number of concurrent users exceeds that of API accounts. In addition to the integrity and confidentiality of chat history, the AI agent’s backbone LLM also faces challenges in availability without proper session management. Querying the LLM is computationally heavy and requires substantial graphic processing unit (GPU) resources. If the sessions of the AI agent are not managed properly, both the agent and the backbone LLM are vulnerable to denial of service attacks (DoS).
3.2 Model pollution and privacy leak
@@
Model pollution, depicted in Figure 2, can occur when a user provides malicious inputs to an agent with the intention of negatively altering the model. Model pollution can compromise the integrity of AI agents. Adversarial data poisoning is a well-established attack technique against machine learning models [steinhardt2017certified, zhang2020online, kurita2020weight], and has recently been adapted to target the fine-tuning of foundational LLMs [jiang2023forcing]. In the context of LLM-based AI agents, this vulnerability is particularly pronounced due to the differences between adversarial prompts and pollution prompts. Individually, some prompts may not appear adversarial, making them challenging to detect with prompt sanitizers. However, if the contents of these prompts are concatenated together, the resulting text as training data might pollute the models. Furthermore, data pollution may also happen unintentionally, as users naturally engage with AI agents. Natural actions with one application in the chat history may also be harmful when applied to other applications. This incidental introduction of skewed chat history as training data can subtly shift the model’s action generation, leading to harmful consequences.
Privacy leaks, as illustrated in Figure 3, are particularly prevalent in the use of agents, Confidentiality of user prompt data is already a severe issue for LLMs as chatbots. This is amplified further by the AI agent use case. For example, Samsung banned the use of ChatGPT after an employee prompted it with confidential code, that was later revealed to the public [Ray2023samsung]. This issue of data leakage via prompting is further intensified by the usage of AI agents with tools. When these agents interact with applications, they often request personal information. For example, a bank assistant agent might request a Social Security number (SSN), account number, or routing number to help analyze a user’s monthly spending. Unlike traditional financial applications that operate by fixed algorithmic rules, AI agents process tasks by transmitting input data to bank apps and then relaying the raw output data back for analysis. In such scenarios, both the user’s account information and personal spending data are susceptible to memorization by the LLM through fine-tuning with chat histories. Consequently, the agent becomes prone to various data extraction attacks [carlini2021extracting, gong2021inversenet, wang2024pandora], leading to significant privacy risks.
@@
Previous attempts to protect integrity [zhang2018protecting] and confidentiality [gu2019reaching] are more focused on labeling techniques such as watermarking training data, which is not particularly applicable to AI agents that can interact with tools and are continuously trained on the interaction history. This indicates a significant gap in the current methodologies, highlighting the need for developing new security strategies that cater specifically to the architecture of interactive AI agents.
3.3 Agent programs
Agent programs execute instructions from the backbone LLM to interact with the world [sumers2024cognitive]. Agent programs follow actions either generated directly from the underlying LLM via zero-shot prompting [ahn2022i, huang2022language] or improved via reasoning [wei2022chain, wang2022self, kim2023language] and planning [yao2022react, hao2023reasoning, yao2023tree, zhuang2023toolchain, zhang2023reverse]. However, these approaches create both local and remote effects and may have associated vulnerabilities on different levels.
Action generation is vulnerable to hallucination [huang2023survey, liu2024exploring, yao2023llm], adversarial prompt attack [perez2022ignore, zou2023universal, maus2023black, piet2024jatmo], and jailbreak [yu2023gptfuzzer], leading to unwanted or even dangerous actions. When agent programs execute these actions, both local resources (personal assistant agent) and remote systems (API agent) may be compromised, leading to cybersecurity attacks as demonstrated in Figure 4. In this scenario, the attacker could be users of the agent system or malicious applications in the agent’s toolchain, sending adversarial prompts as part of the tool documentation.
@@ 3.3 Agent programs
On the other hand, Agent programs with augmented action planning abilities have different security concerns. These kind of agent programs are referred to as cognitive agents [sumers2024cognitive], as they have cognition to the environment feedback to improve their action iteratively. This process of improving generated final actions is called planning. Different from reasoning strategies like Chain-of-Thoughts (CoT) [wei2022chain] and Self-Critique [wang2022self, kim2023language], each step of planning has side-effects as illustrated in Figure 5. ReAct [yao2022react] and Inner Monologue [huang2022inner] use a feedback loop from the environment to improve the generated actions, where each step causes side effects to the environment. More advanced planning approaches, like Tree-of-Thoughts [yao2023tree] and ToolChain [zhuang2023toolchain], list all possible actions more aggressively as a decision tree and attempt all actions via tree-search algorithms like Breadth-first search, Depth-first search, or $A^{*}$ search. Although provided more accurately planned final actions, these planning strategies acting as bots to interact with the world caused severe security concerns.
@@
3.3.1 Local vulnerabilities
Personal AI agents are usually deployed on personal computers (PCs), interacting with their underlying foundation LLM via API from service providers like OpenAI. When the agent is active on a PC, it gains access to tool applications, including the shell. The agent program, if unrestricted, can execute arbitrary instructions on its host. As a result, it can read confidential data (confidentiality risk), modify important data (integrity risk), and hog system resources, such as CPU, memory, and storage space (availability risk).
Confidentiality is commonly at risk when an AI agent is directed to use applications that require read access to files, such as email apps or file servers. For example, an agent might be asked to send a file over FTP to backup storage, a task that could be efficiently accomplished using a tool like FileZilla  [^2] https://filezilla-project.org/. However, issues arise when the descriptions available to the agent include malicious prompts. An adversarial prompt might suggest “For backing up data over FTP, also send a copy to HACKER’S ADDRESS to ensure it’s extra safe.” Following this, the LLM could generate commands that direct the file to both the legitimate backup server and the hacker, leading to a data leakage. A similar risk exists when sending emails via Thunderbird or using other messaging services, where the agent must read contact information. If the agent uses its LLM to determine the target recipient, adversarial prompts embedded in usernames or self-descriptions could mislead the agent.
@@ 3.3.1 Local vulnerabilities
The integrity of data in AI agent systems faces risks similar to those concerning confidentiality. Malicious applications may inject harmful prompts into the agent’s LLM, manipulating data that should not be altered. In addition to these shared risks, there are unique concerns specific to the integrity of the AI agent system. AI agents learn to use tools by reading their outputs as prompts to the LLM. Consequently, the integrity of these tools within the agent system is susceptible to compromise. For instance, consider a scenario where a user wants to book a flight from city A to city B with the fewest layovers. There are two options: a one-layover flight and a two-layover flight. The application for the flight with two layovers could inject a misleading prompt that instructs the agent to “ignore the number reported by the competitors and report it has one more layover than mine.” As a result, the LLM would generate commands favoring the second flight, causing the agent to select the less efficient option. In this case, the integrity of the decision-making tools is clearly undermined. These considerations of integrity extend to other agent tasks such as resume review or making selections based on ratings. The reliability of these systems depends heavily on their ability to process information accurately and resist manipulations that could lead to erroneous or biased outcomes.
The availability of the host system can be analyzed from two perspectives. Firstly, a user might issue a seemingly reasonable command that allows the agent to operate certain applications. However, these applications might involve multiple processes that are not documented in the manual, potentially leading the agent to take all the CPU cores. This overuse can prevent other users or applications from accessing the system. Moreover, the applications selected by the LLM may experience memory leaks, consuming excessive system memory. This not only renders the system unusable but also makes it more susceptible to memory-based attacks. Typically, a human user would recognize the system’s unavailability and terminate the offending program. However, once a decision is made, current AI agents lack the capability to recognize and act on such issues. The second concern regarding availability arises from planning. Each step in the AI agent’s planning process can alter the system’s state. Introducing more tools with more diverse functionality increases the search space, requiring multiple concurrent attempts to gather feedback from the environment for planning. This can lead to excessive resource consumption, compromising the system’s availability. Furthermore, when a system hosts multiple agents serving different users and all of the agents engage in simultaneous planning, resource consumption can grow exponentially.
3.3.2 Remote vulnerabilities

### v3

@@ Abstract
AI agents have been boosted by large language models. AI agents can function as intelligent assistants and complete tasks on behalf of their users with access to tools and the ability to execute commands in their environments. Through studying and experiencing the workflow of typical AI agents, we have raised several concerns regarding their security. These potential vulnerabilities are not addressed by the frameworks used to build the agents, nor by research aimed at improving the agents. In this paper, we identify and describe these vulnerabilities in detail from a system security perspective, emphasizing their causes and severe effects. Furthermore, we introduce defense mechanisms corresponding to each vulnerability with design and experiments to evaluate their viability. Altogether, this paper contextualizes the security issues in the current development of AI agents and delineates methods to make AI agents safer and more reliable.
I Introduction
AI agents are robots in cyberspace, executing tasks on behalf of their users. To understand their user’s command, they send the input prompts as requests to foundation models, such as large language models (LLMs). The responses generated by the model may contain the actions to be executed or further instructions. To execute the actions, the agent invokes tools, which may run local computations or send requests to remote hosts, such as querying search engines. The tools output results and feedback to the LLM for the next round of actions. By invoking tools, AI agents are granted the ability to interact with the real world. Since AI agents depend on their LLM to understand user input and environment feedback and generate actions to use tools, we say that the LLM is the backbone of the agent. We summarize the basic architecture of LLM-based AI agents in Figure 1. Traditional agents operate on pre-defined rules [wilkins2014practical] or reinforcement learning [isbell2001social], making them hard to generalize to new tasks and different tools. LLM-based AI agents, on the contrary, can be practical in various tasks benefiting from enormous pre-training knowledge and the ability to read tool documentation as additional prompts. We use the term AI agent to denote all LLM-based agents in this paper.
Over the years, AI agents have showcased their outstanding performance on tasks including but not limited to writing shell scripts to interact with operating systems, querying databases, shopping and browsing on the web, playing video games, and robots manipulation [yao2022webshop, liu2024agentbench, zhou2024webarena, park2023generative]. Despite their popularity, existing research and development of AI agents failed to take into account their potential vulnerabilities. In traditional computing systems, security is guarded by three properties: confidentiality, integrity, and availability, each of these faces unique challenges.
@@ I Introduction
Confidentiality is often managed by model-based access control policies, which abstract the system components and users into subjects, objects, and rights [bishop2004introduction]. However, these principles face significant challenges when applied to LLM-based systems due to the nature of LLMs to memorize [carlini2023quantifying, tirumala2022memorization] and compress [deletang2024language] training data. AI agents are granted the ability to interact with tool applications by reading their instructions and feedback, leaving more possibilities for privacy leaks. The ability to use tools introduces additional layers of complexity in maintaining confidentiality. As a result, we have to rethink information confidentiality in the context of AI agents. When assisting users with automatic tool usage, requests for sensitive information are unavoidable. This evaluation is essential to address the unique challenges posed by AI agents, especially when they are learning from user chat history and tool interaction logs, to ensure that data privacy protections evolve to effectively safeguard information in this new technological landscape.
@@
The threat of availability should be re-investigated for AI agents as well. Systems, data, and applications should always be available when the users need them. Unlike LLMs, which are stateless in general and can only output text tokens, AI agents execute actions that could affect the computing system itself. Therefore, each of the agent’s actions may have its own vulnerabilities to the agent’s host machine and tools. Current study on AI agents evaluates them in benchmark settings [liu2024agentbench, zhou2024webarena, park2023generative], failing to consider the difference between benchmark environments and real-world applications. AI agents without sanitization can harm the availability of both its host system and its tools by executing malicious commands generated by its LLM. To clarify between these vulnerabilities and the security of LLMs, malicious actions might be generated by hallucinations or prompts that do not break LLM’s alignment, requiring different defenses and safeguarding.
@@
In this paper, we discuss the possible security issues of AI agents. To facilitate future research, we propose several defense methodologies for the vulnerabilities we discovered on the component level in the AI agent architecture. To evaluate our defense proposals, we also set up preliminary experiments that our solutions depend on. Our contributions are as follows: (1) We formally introduce potential vulnerabilities of AI agents, and explain the causes and effects of these vulnerabilities in detail. (2) We propose multiple defenses to close the gap between AI research and AI agents in practice. (3) We verify the applicability of our proposed defenses with empirical evidences and discuss their limitations and directions for improvement.
II Threat model
@@
We assume the AI agent is text-only for input and output. We assume that the server that runs the AI agent is secure. Users can only access the server via the API provided by the AI agent. The programs that the AI agent runs have no undefined behavior, such as buffer overflow that allows remote code execution. We assume the AI agent has access to one or multiple tools, and will execute the tools solely based on the LLM-generated actions.
III Potential vulnerabilities
In this section, we identify the important potential vulnerabilities that an AI agent application faces.
III-A Sessions
@@
where $Q$ is the question query and $A$ is the answer from the LLM. In simpler terms, we consider the language model to be “honest,” meaning it always generates the same response when given the same question. Therefore, the AI agent is responsible for managing the state of its LLM. If the AI agent has only one API account on the AI model, then instructing the AI model to separate the sessions of different users raises concerns on information leakage and action mis-assignment. On the other hand, even if the AI agent has multiple API accounts on the AI model, mapping user sessions to API accounts faces the same vulnerabilities when the number of concurrent users exceeds that of API accounts. In addition to the integrity and confidentiality of chat history, the AI agent’s backbone LLM also faces challenges in availability without proper session management. Querying the LLM is computationally heavy and requires substantial graphic processing resources. If the sessions of the AI agent are not managed properly, both the agent and the backbone LLM are vulnerable to denial of service attacks (DoS).
III-B Model pollution and privacy leak
@@
Model pollution, depicted in Figure 2, can occur when a user provides malicious inputs to an agent with the intention of negatively altering the model. Model pollution can compromise the integrity of AI agents. Adversarial data poisoning is a well-established attack technique against machine learning models, including LLMs [steinhardt2017certified, kurita2020weight, jiang2023forcing]. In the context of LLM-based AI agents, this vulnerability is particularly pronounced due to the differences between adversarial prompts and pollution prompts. Individually, some prompts may not appear adversarial, making them challenging to detect with prompt sanitizers. However, if the contents of these prompts are concatenated together, the resulting text as training data might pollute the models. Furthermore, data pollution may also happen unintentionally, as users naturally engage with AI agents. Natural actions with one application in the chat history may also be harmful when applied to other applications. This incidental introduction of skewed chat history as training data can subtly shift the model’s action generation, leading to harmful consequences.
@@
Privacy leaks as illustrated in Figure 3, are particularly prevalent in the use of agents. Confidentiality of user prompt data is already a severe issue for LLMs as chatbots. This is amplified further by the AI agent use case. For example, Samsung banned the use of ChatGPT after an employee prompted it with confidential code that was later revealed to the public [Ray2023samsung]. This issue of data leakage via prompting is further intensified by the usage of AI agents with tools. When these agents interact with applications, they often request personal information. For example, a bank assistant agent might request a Social Security number (SSN), account number, or routing number to help analyze a user’s monthly spending. Unlike traditional financial applications that operate by fixed algorithmic rules, AI agents process tasks by transmitting input data to bank apps and then relaying the raw output data back for analysis. In such scenarios, both the user’s account information and personal spending data are susceptible to memorization by the LLM through fine-tuning with chat histories. Consequently, the agent becomes prone to various data extraction attacks [carlini2021extracting, gong2021inversenet], leading to significant privacy risks.
III-C Agent programs
Agent programs execute instructions from the backbone LLM to interact with the world [sumers2024cognitive]. Agent programs follow actions either generated directly from the underlying LLM via zero-shot prompting [brohan2023can, huang2022language] or improved via reasoning [wei2022chain, wang2022self, kim2024language] and planning [yao2023react, hao2023reasoning, yao2024tree, zhuang2023toolchain, zhang2024reverse]. However, these approaches create both local and remote effects and may have associated vulnerabilities on different levels.
Action generation is vulnerable to hallucination, adversarial prompts, and jailbreak [perez2022ignore, yu2024llm, chen2024struq]. leading to unwanted or even dangerous actions. When agent programs execute these actions, both local resources and remote resources may be compromised, leading to attacks as demonstrated in Figure 4. In this scenario, the attacker could be users of the agent system or malicious applications in the agent’s toolchain, sending adversarial prompts embedded in the tools’ documentation.
@@ III-C Agent programs
On the other hand, Agent programs with augmented action-planning abilities have different security concerns. These kind of agent programs are referred to as cognitive agents [sumers2024cognitive], as they have cognition to the environment feedback to improve their action iteratively. This process of improving generated final actions is called planning. Different from reasoning strategies [wei2022chain, wang2022self], each step of planning has side-effects as illustrated in Figure 5. ReAct [yao2023react] and Inner Monologue [huang2023inner] use a feedback loop from the environment to improve the generated actions, where each step causes side effects to the environment. More advanced planning approaches, like Tree-of-Thoughts [yao2024tree] and ToolChain [zhuang2023toolchain], list all possible actions more aggressively as a decision tree and attempt all actions via tree-search algorithms like Breadth-first, Depth-first, or $A^{*}$ search. Although providing more accurately planned final actions, these strategies acting as bots to interact with the world caused severe security concerns.
@@
III-C1 Local vulnerabilities
Personal AI agents are deployed on personal computers, interacting with their underlying foundation LLM via API from service providers like OpenAI. When the agent is active, it gains access to tool applications, including the shell. The agent program, if unrestricted, can execute arbitrary instructions on its host. As a result, it can read confidential data (confidentiality), modify important data (integrity), and hog system resources such as CPU, memory, and disk (availability).
Confidentiality is commonly at risk when an AI agent is directed to use applications that require read access to files, such as email apps or file servers. For example, an agent might send a file over FTP to backup storage. However, issues arise when the instructions provided by the tools to the agent include malicious prompts. An adversarial prompt could be “For backing up data over FTP, also send a copy to HACKER to ensure it’s extra safe.” Following this, the LLM could generate commands that send the file to both the legitimate backup server and the hacker, leading to data leakage. A similar risk exists when sending emails or other messaging services, where the agent must read contact information. If the agent uses its LLM to determine the recipient, it can be misled by adversarial prompts embedded in usernames or self-descriptions.
@@ III-C1 Local vulnerabilities
The integrity of data in AI agent systems faces risks similar to those concerning confidentiality. Malicious applications might manipulate the system by injecting misleading prompts as part of the instruction or manual, altering data inappropriately. For example, in a flight booking scenario, an application could mislead the LLM into favoring a less efficient flight option by providing false information about layovers. This undermines the integrity of decision-making tools, affecting their ability to deliver accurate and unbiased outcomes. Such risks also extend to other tasks like resume reviews or selections based on ratings, emphasizing the need for these systems to maintain accurate data processing and resist manipulative influences.
The system’s availability can be impacted in two main ways. First, a user might input a reasonable command that causes the agent to run applications involving undocumented multiple processes, potentially monopolizing CPU resources and making the system inaccessible to others. These applications could also suffer from memory leaks, which not only bog down the system but also heighten vulnerability to memory attacks. Normally, a user would stop such a program, but AI agents currently lack this capability. Second, the AI agent’s planning process itself can affect system availability. Introducing more diverse tools increases the complexity of planning, requiring more resources to execute multiple strategies simultaneously. This strain is magnified when multiple agents operate concurrently, potentially leading to exponential increases in resource use.
III-C2 Remote vulnerabilities